OpenStudy (anonymous):

In 1960, the life expectancy in the UK was 71.1 years. In 2008, it was 79.9 years. Let F(t) = life expectancy and t = the number of years since 1960. a) Find a linear function that fits the data. b) Use the function of Part (a) to estimate the life expectancy in 2010. Show your work not just the final answer.

OpenStudy (jamesj):

You just top posted, but you're not the only one. I gave you all the clues you needed. Post what I wrote again and I'll help you more.

OpenStudy (jamesj):

So ... a linear function is of the form f(x) = mx + c. Draw a graph with the data in this problem: you have two points (x,y) = (1960, 71.1) and (x,y) = (2008,79.9) Now get a ruler and draw the straight line through these points. What is the slope of this curve m? Now solve for c. Can you do that much?

OpenStudy (anonymous):

lol yes i can thank you

OpenStudy (jamesj):

Tell me what values of m and c you get.

OpenStudy (anonymous):

i got 8.8 and 48

OpenStudy (anonymous):

i did y2-y1 over x2-x1

OpenStudy (jamesj):

That's correct in principle, but you have made some sort of error (y2-y1)/(x2-x1) = (79.9 - 71.1)/(2008-1960) = .... NOT 8.8 but ...

OpenStudy (anonymous):

9?

OpenStudy (jamesj):

Seriously? Cmon. Do the math.

OpenStudy (jamesj):

m = (y2-y1)/(x2-x1) = (79.9 - 71.1)/(2008-1960) = 8.8/48 = 0.1833. Now, solve for c.

OpenStudy (anonymous):

oh i c. I was skipping a step...but i do not know hwo to solve for c, that is why i am here

OpenStudy (jamesj):

f(x) = mx + c. So c = f(x) - mx for all x. Take x = 1960. Then f(x) = 71.1 and hence c = f(x) - mx = ....

OpenStudy (anonymous):

if i am adding i got 2031

OpenStudy (anonymous):

or 1888

OpenStudy (jamesj):

Look at the formula and substitute: c = f(x) - mx = 71.1 - (0.1833)x(1960) = 71.1 - 359.3 = -288.3 This means you now know the linear function f(x) = mx + c, as you know both m and c. Clear?